OpenStudy (anonymous):

Convert the equation to polar form. 2x=2y

OpenStudy (anonymous):

x=rcos(tehta) y=rsin(theta) 2x=2y divide both sides by 2 and get: x=y thus : Cos(theta)=sin(theta)

OpenStudy (anonymous):

\[\theta=\pi/4 + 2n\pi\]

OpenStudy (anonymous):

are those in polar form?

OpenStudy (anonymous):

it goes in that direction